1. Blue taxi inc. charges $7.50 for the first 4 miles and 1 dollar for each additional miles. The table shows the cost of travelling from 4 to 10 miles.

Mileage Cost

4. 7.50

5. 8.50

6. 9.50

7. 10.50

8. 11. 50

9. 12.50.

10. 13.50

  • Write the recursive formula.
  • Write the explicit formula


2. If the sequence formula is given as tn = n2 -1, n less than or equal to 1.

  • Write the first five terms of the sequence
  • Write the 55th term
1
Expert's answer
2021-02-24T12:13:35-0500

1.Denote by cn the cost of travelling for n miles. Then

cn=cn-1+1 if n>4 and cn=cn-1=7.5 else.

Explicit formula:

cn=n+3.5 if n>4 and cn=7.5 else.


2. tn = n2 -1, n less than or equal to 1.

t1 = 12-1 = 0

t2 = 22-1=3

t3 = 32-1=8

t4 = 42-1=15

t5 = 52-1=24

t55 = 552-1=3024